In our simulation of sexual selection, what is the allele frequency of "B" and "b" from generation 1 to 3?
what is the trend of each allele?
what do you think causes the change?

4 in le Pe male 12 3 4 38 21 0 5 3 bb 4 3 3 y 10 166 Exercise 12 jEvolution: Lab Report
Answer #1

Gen 1:

  • frequency of B = 14/40 = 0.35 ; frequency of b = 26/40 = 0.65

Gen 2:

  • B = 6/22 = 0.27 ; b = 16/22 = 0.73

Gen 3:

  • B = 3/24 = 0.125 ; b = 21/24 = 0.875

Frequency of allele B is decreasing in the population while that of allele b is increasing.

One of the reasons for the trend could be that allele B is recessive and allele b dominant.